OpenStudy (anonymous):

z+(-3.19)=5.83 what is the value of z

OpenStudy (farmdawgnation):

So it may help to rewrite this like so: z - 3.19 = 5.83 Does that help?

OpenStudy (farmdawgnation):

If not I (or someone else) would be more than happy to help walk you through it. :)

OpenStudy (anonymous):

yes but i need the answer

OpenStudy (farmdawgnation):

Ok, but I can actually teach you how to do it if you like. That way you'll be able to do it yourself. Or, we can just give you the answer. I'd prefer to help teach you so you don't have to keep asking.

OpenStudy (anonymous):

ok...

OpenStudy (farmdawgnation):

Ok, so essentially, you just need to do some things to both sides of the equality to reveal what Z actually stand for. So, we have for this problem: z - 3.19 = 5.83 Now, really, we need to get rid of this nasty -3.19 and z will be left all by its lonesome and give us our answer. So, to get rid of -3.19 in this problem we can add 3.19 to each side. So, -3.19 + 3.19 = 0. And 5.83 + 3.19 = 9.02 That means that we're left with: z = 9.02 Make sense?

OpenStudy (anonymous):

si

OpenStudy (anonymous):

gracias senor

OpenStudy (farmdawgnation):

Glad I could help!